The constant of a spring is 480 N/m.


How much elastic potential energy is stored in the spring if the spring is compressed a distance of 6.4 cm?


What is the force being used to compress the spring?


1
Expert's answer
2021-07-05T17:59:49-0400

The potential energy is given as follows:


"W = \dfrac{kx^2}{2}"

where "k = 480N\/m" and "x = 6.4cm = 0.064m". Thus, obtain:


"W = \dfrac{480\cdot 0.064^2}{2} \approx 0.98J"

The force is given by the Hook's law:


"F = kx = 480\cdot 0.064 \approx 31N"

Answer. 0.98J, 31N.